WHAT IS (W+5)+W+(W+5)+w=82 ????????????
PLEASE AWNSER

Answer by Alan3354(69443)

WHAT IS (W+5)+W+(W+5)+w=82 ?
---------
It's an equation in w
-----------
-----------
If you want to solve for w, you should say, "Solve for w.", not "what is..."
-------
(W+5)+W+(W+5)+w=82
Assuming W = w:
(W+5)+W+(W+5)+w=82
4w + 10 = 82
4w = 72
w = 18
-------